### v1

# Nonparametric regression for cost-effectiveness analyses with observational data - a tutorial

Abstract
Healthcare decision-making often requires selecting among treatment options under budget constraints, particularly when one option is more effective but also more costly. Cost-effectiveness analysis (CEA) provides a framework for evaluating whether the health benefits of a treatment justify its additional costs. A key component of CEA is the estimation of treatment effects on both health outcomes and costs, which becomes challenging when using observational data, due to potential confounding. While advanced causal inference methods exist for use in such circumstances, their adoption in CEAs remains limited, with many studies relying on overly simplistic methods such as linear regression or propensity score matching. We believe that this is mainly due to health economists being generally unfamiliar with superior methodology. In this paper, we address this gap by introducing cost-effectiveness researchers to modern nonparametric regression models, with a particular focus on Bayesian Additive Regression Trees (BART). We provide practical guidance on how to implement BART in CEAs, including code examples, and discuss its advantages in producing more robust and credible estimates from observational data.
volume: TBAissue: TBA
[…]
, , and
1 Introduction
A central challenge in healthcare decision-making is determining which interventions to implement or reimburse within the limits of constrained budgets. This often necessitates choosing between competing treatment options for the same medical condition, where one treatment proves to be both more effective and more costly than the other. Cost-effectiveness analyses (CEAs) are designed to inform such choices by assessing whether the additional health gains of a more effective treatment justify its additional costs. A key component of a CEA is the estimation of the treatment effects on both patients’ health outcomes and healthcare costs using available data.
If this data is obtained from a randomized trial, where the treatment allocation does not depend on the patients’ baseline characteristics, the statistical estimation is relatively simple. If the data is observational, and the treatment allocation is hence related to the patients’ baseline characteristics, the estimation procedure is much more complex. Despite the availability of several textbooks on the analysis of observational data (Imbens and Rubin, 2015; Aronow and Miller, 2019; Hernán and Robins, 2024), appropriate methodology is not widely used in the field of CEA. That is, many researchers rely on simple linear regression or propensity score matching, preventing them from capturing complex, nonlinear relationships or interactions within the data. While there are some proposals for more sophisticated statistical methods in CEAs (Kreif et al., 2013; Chen et al., 2024), we do not think they go far enough, since they ultimately still rely on restrictive parametric models.
We see several potential reasons for the over-reliance on such simple models. One is simply the unawareness of useful alternatives, or of the advantages which these alternatives offer. Another is the added complexity of nonparametric models, with respect to both usage and interpretation. In this paper, we therefore aim to enlargen the toolkit available to cost-effectiveness research, by giving an accessible guideline on how to use state-of-the-art nonparametric regression methods for CEAs. Our focus will be on Bayesian Additive Regression Trees (BART), which has recently been extended to work in the cost-effectiveness setting.
2 Preliminaries
We begin by explaining what researchers aim to learn from CEAs using observational data. Our presentation will be quite informal; a more rigorous version may be found in Esser et al. (2024). We shall not consider more general issues in the analysis of observational data, such as well-defined patient populations, treatment options, follow-up periods and so on. Such aspects are very important, arguably more important than the statistical aspects, but they apply to CEAs in exactly the same way as they do to more general epidemiological analyses. We thus refer the reader to canonical sources of these issues, such as Hernán and Robins (2016).
In a typical CEA, we are faced with two treatment options, $t=0$ and $t=1$. Often the former is called the control, and the latter the intervention. The fundamental goal is to estimate, given available data, the probability that the intervention is cost-effective. To make this possible, we consider two essential quantities:
  • $\Delta_{c}$, the average treatment effect on the costs. Put simply, $\Delta_{c}$ tells us how much more we would need to pay per patient, on average, if every patient received the intervention instead of the control. It is typically expressed in the local currency which healthcare authorities are interested in — for instance, the Dutch National Health Care Institute uses EUR (Hakkaart-van Roijen et al., 2015).
  • $\Delta_{q}$, the average treatment effect on the patient’s health [^1] Instead of health many authors use the term effect here; see for example Drummond et al. (2015). We avoid this terminology, since we consider it confusing: $\Delta_{q}$ would then be the ”effect on the effect”.. This quantity tells us how much healthier the patients would be, on average, if every patient received the intervention instead of the control. The patients’ health can be measured in many different ways. Some are disease-specific, others are intentionally non-specific, in order to make the health state of patients comparable across different diseases. Of the latter the most commonly used is the Quality-adjusted life year (QALY) (Drummond et al., 2015).
In health economic research, we frequently encounter the situation that both treatment effects are positive, which means that the intervention is both more effective and more expensive than the intervention. A natural question then arises: do the added health gains of the intervention justify the increase in costs? To answer this, we need to quantify how much we are willing to pay for an additional unit of health. This quantify is called the willingness-to-pay (WTP) and is customarily assigned the symbol $\lambda$.
Example 1. Suppose that costs are measured in EUR, and $\Delta_{c}=1000$; health is measured in QALYs, and $\Delta_{q}=0.1$. Then, by implementing the intervention, we gain an average health benefit of $0.1$ QALYs per patient, while, having to pay an additional $1000$ EUR. We hence need to pay $10000$ EUR in exchange for one QALY gained (because $1000/0.1=10000$). The intervention is cost-effective if and only if $\lambda$ is greater than $10000$. In other words, the intervention is cost-effective if and only if $\lambda*0.1-1000>0$.
The decision process in the preceding example can be formalised by working with the the incremental net benefit (INB):
$\text{INB}_{\lambda}\coloneqq\lambda\Delta_{q}-\Delta_{c}.$
This formula helps us decide if the extra health we get from the intervention is worth the extra cost. We do this by putting a monetary value on health ($\lambda$) and checking if the benefits of the intervention outweigh its additional costs. More precisely: if the INB is zero or negative, we conclude that the intervention is not cost-effective and should hence not be implemented. Conversely, if the INB is positive, we consider the intervention to be cost-effective and worthy of implementation.
In most applications, the true values of $\Delta_{c}$ and $\Delta_{q}$ are unknown, and so is the true value of $\text{INB}_{\lambda}$ for any given $\lambda$. They thus must be estimated from available data, and there will inevitably be uncertainty about these estimates. In this paper, we adopt the Bayesian point of view when conceptualizing this uncertainty, which means that we treat all unknown quantities as random variables. This allows us to speak about the probability of cost-effectiveness [^2] In the frequentist or ”classical” paradigm, this is not possible, and we would instead have to argue in terms of p-values. See Löthgren and Zethraeus (2000)., which is given by $\Pr(\text{INB}_{\lambda}>0)$ (the probability that, given a specific $\lambda$ value, the INB is positive).
While there are many different methods to estimate $\Delta_{c}$ and $\Delta_{q}$, they are all ultimately based on empirical observations of patient’s cost and health. These observations are often positively or negatively correlated, as illustrated by the following examples.
  • Positive correlation: Stargardt et al. (2014) found that, for patients suffering from acute myocardial infarction, higher use of healthcare resources is associated with higher survival rates.
  • Negative correlation: (Domínguez-Ortega et al., 2015) found that, in asthma care, increased disease severity (hence worse health) is associated with higher healthcare costs.
A proper statistical analysis should account for this correlation, and doing so will also induce correlation in the distribution of $\Delta_{c}$ and $\Delta_{q}$.
Example 2. We now consider a situation where $\Delta_{c}$ and $\Delta_{q}$ are estimated from available data, and we are hence uncertain about their values. This uncertainty is expressed through a probability distribution; an example of which is illustrated in Figure 1. The left column shows two bivariate density plots of $\Delta_{q}$ and $\Delta_{c}$, while the right column shows the probability of cost-effectiveness as a function of $\lambda$. This plot is called a cost-effectiveness acceptability curve (CEAC). While the mean and variance of $\Delta_{c}$ and $\Delta_{q}$ are the same in both rows, the correlation between them differs: it is negative in the top row and positive in the bottom row. This difference in correlation leads to a noticeable change in the CEACs and, consequently, in how the results are interpreted. For example, at a willingness-to-pay of $50000\text{€}$ — a commonly used European WTP for a QALY — the probability of cost-effectiveness is approximately $0.7$ in the top row, but nearly $1$ in the bottom row.
Figure 1: Correlation between $\Delta_{c}$ and $\Delta_{q}$ and its impact on the probability of cost-effectiveness.
3 Observational studies and model specification
As established above, the main task in a CEA is to jointly estimate the treatment effects $\Delta_{c}$ and $\Delta_{q}$ from available data. We shall now review some statistical aspects of this process.
If the data are obtained from a randomized controlled trial, model specification is not critical. That is, we do not need to adjust for patients’ baseline characteristics because, by design, they are unrelated to which treatment the patient receives, and hence simply comparing mean costs and effects in the treatment groups will produce good estimates of the treatment effects. If a researcher insists on adjusting anyway, a simple linear model will perform well (Lin, 2013). Without randomization, however, there will generally be certain baseline characteristics, called confounders, which influence both the treatment choice and the outcomes (i.e. cost and health) and hence need to be adjusted for in the analysis. This adjustment necessitates the use of a statistical model, as we need to specify how exactly the confounders influence the outcomes. Even if all relevant confounders are measured, a badly chosen model can potentially lead to severely biased and inconsistent estimates of the treatment effects (Aronow et al., 2025), making model choice a critical matter when analysing observational data.
With that in mind, we will now critically examine some of the commonly used model choices in CEAs. A frequently used model is seemingly unrelated regression (SUR, Zellner (1962); Ben et al. (2023)), which asserts that
$\begin{pmatrix}c_{i}\\ q_{i}\end{pmatrix}$ $=\begin{pmatrix}\boldsymbol{\beta_{1}}^{\top}\mathbf{x_{i}}\\ \boldsymbol{\beta_{2}}^{\top}\mathbf{x_{i}}\end{pmatrix}+\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix},$ $\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix}\sim\text{N}\Biggl{(}\begin{pmatrix}0\\ 0\end{pmatrix},\boldsymbol{\Sigma}\Biggl{)}$ (1)
We model the patients costs $c_{i}$ and health $c_{i}$ as linear functions of a set of covariates $\mathbf{x_{i}}$, which is some transformation of the patient’s baseline characteristics (so it may include, say, quadratic or interaction terms), plus a bivariate-normally distributed distributed error term. Both the linearity and the normality assumption are questionable, and we shall now scrutinize each of them in turn.
The normality assumption in the SUR model is rarely realistic, particularly for the cost outcome, which can only take non-negative values. However, violations of this assumption are not critical: in a landmark paper, White (1980) showed that, as long as certain mild conditions are met, the estimates from linear regression models remain unbiased and consistent [^3] Consistency means, roughly speaking, that with a sufficiently large sample the estimate will be close to the true value., even when the distribution of the error term is misspecified.
Nevertheless, many authors have promoted the use of generalized linear models (GLMs), which allow the specification of non-normal error terms—for example, by modeling costs using a lognormal or gamma distribution (Baio, 2018; Gabrio et al., 2019). However, this approach has a significant drawback: the results can be very sensitive to the choice of the distribution. A valuable simulation study by Briggs et al. (2005) showed that, if a lognormal distribution is assumed for the costs, but the true distribution is gamma, the estimated mean costs are strongly biased. The sample mean on the other hand, which is essentially a linear regression without covariates, performs well regardless of the true distribution. The same patterns will persist in the more observational setting, with adjustment for baseline covariates. The reason is that, for GLMs with non-normal error terms, there are no theoretical guarantees analogous to those in linear models — that is, if the error term in a GLM is not specified correctly, the resulting estimates may be inaccurate even with large sample sizes. In consequence, choosing the wrong distribution for the outcome can lead to both biased and inconsistent treatment effect estimates.
In our view, the more fundamental problem with the SUR model is the assumed linearity. It is well-known in the causal inference literature that violations of this assumption can lead to severe bias in the estimated treatment effects (Dorie et al., 2019), and this extends without change to the cost-effectiveness setting. We illustrate this through a simple simulation experiment in Appendix A. GLMs suffer from analogous problems. For example, a GLM for costs often assumes that the logarithm of expected costs depends lineary on the covariates (Baio, 2018); an assumption just as restrictive as the linearity in the SUR model. More generally, both SUR and GLMs are examples of parametric models, which make strong assumptions about how the covariates are related to the outcomes. Violations of these assumptions lead to distorted estimates of treatment effects (Dorie et al., 2019).
In order to deal with the aforementioned problem, statisticians have developed regression techniques which do not make strong parametric assumptions about the relationship between the covariates and the outcomes. These so called nonparametric regression models [^4] The term ”nonparametric” is somewhat of a misnomer, since such models do in fact have parameters. However, the number of these parameters (or more technically: the dimension of the parameter space) is not chosen in advance, unlike in, say, the basic linear model. Nonetheless, the terminology is widely used. are often much more accurate than linear models or GLMs. Crucially, analogues for the consistency guarantees of linear regression, under similar conditions, also exist for nonparametric regression models with a misspecified normal error term (Kleijn and Van der Vaart, 2006). The added benefits come at a cost, however, as nonparametric regression methods can be somewhat difficult to use and interpret. In particular, unlike in the basic linear model, the treatment effects of interest are no more represented by a single model parameter (more on this later). We thus intend to make the methodology accessible to applied cost-effectiveness researchers, providing sufficient detail for responsible use, but not more.
4 BART for CEAs
We will focus on a specific nonparametric regression model, namely Bayesian additive regression trees (BART), which has been shown to perform well at estimating treatment effects (Dorie et al., 2019). We have recently extended this model to handle both non-linear relationships and the correlation between costs and health outcomes in CEAs. This extended version is called suBART (Esser et al., 2024), and it is especially useful given the challenges discussed in Section 2.
As the name suggests, BART is a Bayesian statistical model and works by aggregating a set of regression trees. We shall now explain briefly what each of these terms mean.
4.1 Regression trees
Regression models can take many forms. They all have in common that they, in some way, relate an outcome variable to a set of covariates. A regression tree does this by diving the covariate space (the set of all possible values of $\mathbf{x}_{i}$) into distinct ”bins”. Then each of these bins is assigned a single number, which is the ”prediction” or the ”expected outcome” for all observations for which the covariates fall into the corresponding bin.
Example 3. An representation of a simple regression tree is shown on the left of Figure 2. We may imagine this tree giving us the predicted QALY for a particular patient, based on their baseline health state (expressed as a number between 0 and 1). Given a specific value for this baseline utility — say $x$ — how do we find the predicted QALY? We start at the top of the tree, and then move down the levels sequentially. At each step, we follow either the left or the right direction, depending on whether $x$ is smaller or larger than the shown threshold value. Once we reach the end of a branch (where the tree no longer splits), we take the given value as the predicted QALY.
We can also use the tree to plot the expected QALY as a function of $x$. This is illustrated in the right of Figure 2. Where a linear model would have given us a line, the regression tree produces a so-called step function.
{forest}
for tree= grow=south, draw, minimum size=3ex, inner sep=3pt, s sep=7mm, l sep=6mm [$x\leq 0.3$, [$0.2$, edge label=node[midway, left, font=]TRUE , circle,] [$x\leq 0.6$, edge label=node[midway, right, font=] FALSE, [$x\leq 0.5$, edge label=node[midway, left, font=], [0.9, edge label=node[midway, left, font=], circle,] [0.4, edge label=node[midway, right, font=], circle,] ] [0.6, edge label=node[midway, right, font=], circle,] ] ]
Figure 2: Regression tree (left) and implied regression function (right).
Regression trees are somewhat analogous to the decision tree models which health economists sometimes use in simulation experiments (see for example Briggs et al. (2006), chapter 2). However, among other differences, a regression tree is not a priori specified by the researcher, but rather ”learned from the data”.
4.2 BART
The basic idea of BART is to express the outcome as a sum of regression trees plus a normally distributed error term [^5] As noted above, such a normally distributed error term does not pose issues even when the conditional distribution of the outcomes is not normal.:
$y_{i}=\sum^{m}_{t=1}f_{i}\left(\mathbf{x}_{i}\right)+\varepsilon_{i},$
where each $f_{i}$ is a regression tree. Thereby the model combines the contributions of many small regression trees, each capturing different patterns in the data. This approach allows for modelling complex nonlinear relationships between covariates and outcome without making strong parametric assumptions. $200$ are frequently used, since this choice has been found to provide good performance while keeping the computational demands of the model fitting manageable. See Chipman et al. (2010); Tan and Roy (2019) for detailed treatments of the original BART model, and in particular the process by which the trees are learned from the data.
4.3 suBART
The acronym suBART stands for seemingly unrelated BART, a name deliberately reminiscent of the SUR model we described above. SUR consists of two linear models which are coupled through correlated error terms. suBART keeps the latter, but replaces the linear models with additive regression trees. More precisely, the model is given by:
$\begin{pmatrix}c_{i}\\ q_{i}\end{pmatrix}$ $=\begin{pmatrix}\sum^{m}_{j=1}f_{j}(\mathbf{x_{i}})\\ \sum^{m}_{j=1}g_{j}(\mathbf{x_{i}})\end{pmatrix}+\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix},$ $\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix}\sim\text{N}\Biggl{(}\begin{pmatrix}0\\ 0\end{pmatrix},\boldsymbol{\Sigma}\Biggl{)},$ (2)
where each $f_{i}$ and $g_{i}$ is a regression tree.
4.4 A note on Bayesian estimation and MCMC
As the name suggests, BART is a Bayesian statistical method, which means that the model parameters — and functions of these parameters, such as $\Delta_{c}$ — are treated as random variables, rather than unknown constants. We learn about these parameters by examining their posterior distribution, which quantifies how different values of the parameters are more or less likely, taking into account both prior beliefs and the evidence provided by the data.
Bayesian estimation has a long tradition in CEA, with some authors arguing it is the most natural framework to perform cost-effectiveness research under (Baio, 2018). The main reason is that cost-effectiveness researchers want to speak about the probability of cost-effectiveness which is nothing other than $\Pr(\lambda\Delta_{q}-\Delta_{c}>0)$. In other words, we want to make probabilistic statements about our parameters of interest, and this – strictly speaking — is only possible under the Bayesian paradigm.
In addition, there are pragmatic considerations which motivate our use of Bayesian estimation: it enables the use of complex statistical models which would not be usable otherwise, and which perform well even with respect to non-Bayesian criteria. BART is a prime example of this, as shown through the results of the American Causal Inference Conference (ACIC) competition (Dorie et al., 2019).
In practice, Bayesian estimation works by taking a large number of random draws from the posterior distribution. We then use these draws to approximate our quantities of interest, such as the aforementioned probability of cost-effectiveness or confidence intervals for the treatment effects. This approximation will be quite accurate if the number of draws is sufficiently large. One popular procedure to obtain these draws is is called Markov chain Monte Carlo (MCMC), and the collection of draws is called an MCMC sample. The use of MCMC samples is quite similar to that of bootstrap samples, even though the underlying theory is different. For example, to construct a 90% confidence interval, we can simply take the 5th and 95th percentiles of the MCMC sample. Also analogously to bootstrapping, there are no strict guidelines for what the aforementioned ”large number” of draws should be. The usual choice is a few thousand, and we will use $4000$ in the example below.
Details on the specific MCMC method for the suBART model may be found in Esser et al. (2024).
4.5 ps-suBART
In the causal inference literature, one frequently encounters the so-called propensity score (PS), which is the probability that patient received the treatment, given their baseline characteristics. There PS It summarizes all the confounding variables into a single number. The inclusion of the PS as an additional covariate is very beneficial when estimating treatment effects through BART (Hahn et al., 2020; Dorie et al., 2019). In practice, we first estimate the PS for every patient in the data — for example, a BART model that predicts treatment assignment. We then add this PS to the covariates when fitting the BART model for the outcomes. This idea extends in the naturally to suBART, and the resulting ps-suBART model has been shown to perform well in the setting observational CEAs (Esser et al., 2024).
5 Guided example
In this section we guide the reader through the use of suBART. We use a simulated dataset, which is intended as an example and not meant to represent any specific real-world research endeavour. In the supplementary material [^6] See https://github.com/Jonas-Esser/nonparametric_CEA, we provide the full dataset, together will all of the following code. The dataset contains the treatment indicator (which indicates which treatment a patient is part of), the two outcome variables (costs and health), and multiple confounding variables. FOr illustration, we show part of the data in Table 1.
Table 1: Example data.
c q t age sex education
2083.03 0.5865 0 56 0 3
2882.69 0.9512 0 70 1 3
2275.42 0.9149 0 71 1 3
1964.08 0.6929 0 61 0 3
2524.98 0.9133 0 75 1 1
2683.89 0.6261 1 50 1 3
1815.82 0.4071 0 43 1 2
1744.14 0.2123 1 21 1 3
2011.14 0.5122 1 28 0 3
Some observations in the outcome variables are missing. The suBART method can handle these automatically, provided we assume that the data are missing at random —that is, the probability of a value being missing does not depend on the value itself, but may depend on other observed variables (such as confounders).
In order for the following code to work correctly, it is important that the data are formatted properly. Unordered categorical variables with more than two levels (i.e., variables that represent categories without any natural order, such as blood type, region, or treatment center) should be declared as factors. In the simulated dataset, this applies to the trauma_type variable. All other variables should be declared as numbers. Furthermore, all missing observations should be properly formatted as NA values, not as text strings or numbers such as $-99$, which is sometimes done.
1data\$education <- as.factor(data\$education)
2data[data == -99] <- NA
Listing 1: Data formatting
5.1 Data preparation
We put all the covariates into a data frame called X, including the treatment indicator but not the outcomes c and q. The X object should contain only those variables which are supposed to enter into the regression model, and no others. Likewise, we put the outcomes c and q into their own data frame Y.
1X <- data[,c("t", "age", "sex", "education")]
2Y <- data[,c("c", "q")]
Listing 2: Data preparation
5.2 Estimation of propensity scores
We now proceed to estimate the propensity scores (PS). The number of trees and MCMC iterations has been motivated earlier. We also use a so-called burn-in sample of size $1000$; the reasons for this need not concern us here and can be found in any introduction to MCMC.
Once estimated, we append the PS to our covariate data frame, X. For clarity and to keep the original data intact, we store this extended data frame in a new object called X_ps.
1library(subart)
2
3ps_fit <- subart(x_train = X[,c("age", "sex", "education")],
4 y_mat = as.matrix(X\$t),
5 x_test = X[,c("age", "sex", "education")],
6 n_tree = 100,
7 n_mcmc = 2000,
8 n_burn = 1000,
9 varimportance = TRUE)
10
11X_ps <- X
12X_ps\$ps <- apply(pnorm(ps_fit\$y_hat_test), 1, mean)
Listing 3: Estimation of propensity scores
5.3 Further data preparation
As mentioned above, nonlinear models like BART have the property that the treatment effects estimates are not represented by a single model parameter, as is the case in the simple linear model without interaction terms. The treatment effects are instead computed based on predictions from the model, as follows: for each patient, we estimate the expected outcomes under both treatment options and calculate then the average difference between these expected outcomes. The procedure is commonly known as g-computation (Hernán and Robins, 2024).
We shall now lay out how this is done. The idea is to compute the expected outcomes for each patient under both treatment options, and then average the difference between these expected outcomes. More precisely, our strategy may be outlined as follows:
  1. 1.
    Estimate the model parameters.
  2. 2.
    For each patient $i$ predict, according to the model:
    • $\hat{c_{i}}(0)$ and $\hat{q_{i}}(0)$, which are the expected costs and QALY of the patient, had they treatment 00, regardless of what treatment they actually received.
    • $\hat{c_{i}}(0)$ and $\hat{q_{i}}(0)$; defined analogously
  3. 3.
    Obtain $\Delta_{c}$ and $\Delta_{q}$ by averaging the expected values over the sample:
    $\Delta_{c}=\frac{1}{n}\sum^{n}_{i=1}\hat{c_{i}}(1)-\hat{c_{i}}(0)$
    $\Delta_{q}=\frac{1}{n}\sum^{n}_{i=1}\hat{q_{i}}(1)-\hat{q_{i}}(0)$
  4. 4.
    Compute the incremental net benefit for any willingness-to-pay $\lambda$ of interest, through
    $\text{INB}_{\lambda}=\lambda\Delta_{q}-\Delta_{c}$
These steps will be performed for every MCMC draw. For the sake of step $3$, we create another data frame, which contains contains each patient twice: once with the treatment indicator set to 00, and once with it set to $1$. This data frame is called X_test in the code.
1X_test <- rbind(X_ps, X_ps)
2X_test\$t <- c(rep(0, nrow(X)), rep(1, nrow(X)))
Listing 4: Further data preparation
5.4 Fitting the suBART model
We now fit the model, using the previously constructed data frames.
1suBART_ps_fit <- subart(x_train = X_ps,
2 y_mat = as.matrix(Y),
3 x_test = X_test,
4 n_tree = 100,
5 n_mcmc = 2000,
6 n_burn = 1000,
7 varimportance = FALSE)
Listing 5: Fitting suBART
The resulting suBART_fit object contains a variety of objects; we shall only need the entry suBART_fit\$y_hat_test, which are the predictions from Step 5.3.
5.5 Obtaining the results
In this step we apply the calculations from Section 5.3 to all draws in the MCMC sample. We then use the thereby obtained draws of $\Delta_{c}$ and $\Delta_{q}$ to estimate $\Pr\left(\text{INB}_{\lambda}>0\right)$ for a specified range of $\lambda$ values.
1suBART_ps_fit\$ATE <- matrix(NA, nrow = 2000, ncol = 2)
2for (i in 1:2000){
3 suBART_ps_fit\$ATE[i,] <- c(
4 mean(suBART_ps_fit\$y_hat_test[X_test\$t == 1,1,i]) - mean(suBART_ps_fit\$y_hat_test[X_test\$t == 0,1,i]),
5 mean(suBART_ps_fit\$y_hat_test[X_test\$t == 1,2,i]) - mean(suBART_ps_fit\$y_hat_test[X_test\$t == 0,2,i])
6
7 )
8}
9
10MCMC_sample <- data.frame(
11 Delta_c = suBART_ps_fit\$ATE[,1],
12 Delta_q = suBART_ps_fit\$ATE[,2]
13)
14MCMC_sample\$INB20 <- 20000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
15MCMC_sample\$INB50 <- 50000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
16
17lambda <- seq(0, 50000, length.out = 1000)
18p <- rep(NA, length(lambda))
19for (i in 1:length(lambda)){
20 p[i] <- mean(lambda[i]*MCMC_sample\$Delta_q - MCMC_sample\$Delta_c > 0)
21}
22CEAC <- data.frame(lambda, p)
Listing 6: Postprocessing of MCMC output
5.6 Plotting the results
To finish the analysis, we plot the cost-effectiveness plane and the cost-effectiveness acceptability curve (Willan and Briggs, 2006), using the ggplot2 package (Wickham, 2016).
1library(ggplot2)
2
3
4CEplane <- ggplot(MCMC_sample) +
5 geom_point(mapping = aes(Delta_q, Delta_c), size = 1, alpha = 0.3, show.legend = FALSE) +
6 geom_point(aes(mean(Delta_q), mean(Delta_c)), color = "red", size = 2) +
7 geom_vline(xintercept = 0, linetype = "dashed") +
8 geom_hline(yintercept = 0, linetype = "dashed") +
9 labs(x = expression(Delta[q]), y = expression(Delta[c])) +
10 scale_x_continuous(limits = c(-0.05, 0.15)) +
11 scale_y_continuous(limits = c(0, 1000)) +
12 theme_classic() +
13 theme(text = element_text(size = 14))
14
15
16CEAC <- ggplot(data = CEAC) +
17 geom_line(aes(x = lambda, y = p), linewidth = 0.8) +
18 geom_vline(xintercept = 0, linetype = "dashed") +
19 ylim(0,1) +
20 labs(x = expression(lambda)) +
21 labs(y = "Probability of cost-effectiveness") +
22 theme_classic() +
23 theme(text = element_text(size = 14), axis.line.y = element_blank(), legend.position = "right") +
24 scale_x_continuous(label=paste0((0:5) * 10, "k")) +
25 geom_vline(xintercept = 0) +
26 geom_hline(yintercept = 0)
27
28grid.arrange(CEplane, CEAC, nrow = 1)
Listing 7: Plotting the results
Figure 3: Cost-effectiveness plane and CEAC.
6 Discussion
In this tutorial, we illustrated how to perform statistical estimation in cost-effectiveness analysis with observational data.
We focused on the suBART model, both due to personal affinity with it, and because it has been shown to perform well in the aforementioned use case. We stress, however, that our main intention is not a presentation of this model, but rather to provide more applied cost-effectiveness researchers a more general framework for analysing observational data. suBART is merely an example, and there are alternatives which can be used in a similar way, and which we also expect to perform well. In principle, any non-Bayesian method for estimating $\Delta_{c}$ and $\Delta_{q}$ could be used, in conjunction with bootstrapping. The MCMC sample in Section 6 would then be replaced by the bootstrap sample, and the rest of the analysis proceeds in the same way. Such an analysis is no longer Bayesian, and hence the interpretation of the results changes: strictly speaking, we can no longer speak of the ”probability of cost-effectiveness”, and must instead argue in terms of frequentist p-values. See Löthgren and Zethraeus (2000) for details. Also, as mentioned earlier, such approaches may be infeasible in practice, due to computational constraints. Suitable choices for the estimator could include, for example, the Causal Forest of Wager and Athey (2018) or Double Machine Learning (Knaus, 2022).
We now discuss some limitations of our presented approach (which are however shared by the aforementioned alternatives) as well as possible remedies.
  • suBART can only deal with missing observations in the outcome variables, not in the covariates. The technical reason for this is that the model is formulated conditional on the covariates, and hence makes no assumptions about the distribution of the covariates. In principle this is a good thing, because the analysis is therefore not dependent on such assumptions, but it also means that there is no way to impute missing covariates. Conceptually it seems most appealing to directly impute the missing covariates using the BART machinery, as proposed by Xu et al. (2016). While applying this idea to suBART is conceptually straightforward, a computational implementation is currently not available. As a pragmatic alternative, we suggest combining suBART with multiple imputation in the same way as is commonly done when performing CEAs with trial data. The procedure can be sketched as follows: we create multiple complete dataset using standard imputation techniques and compute, for each of these datasets, the point estimates and covariance matrix of $\Delta_{c}$ and $\Delta_{q}$ (see Listing 6). These are then pooled using Rubin’s rules, and the pooled estimates may be used to construct the CEAC. In this context we refer to Harel and Zhou (2007) for a general review of multiple imputation, and Ben et al. (2023) for specific guidance on imputation in trial-based CEAs.
  • We focussed on the basic setting in observational studies, where the treatment for each patient is assigned once and then not changed anymore. Most cost-effectiveness studies in the literature are of this kind, but there is also increasing interest in performing CEAs with so called time-varying treatments [add reference]. In such analyses, the patients may switch from one treatment to another during follow-up. The probability of such switches happening may depend on confounders, and these confounders may themselves change over time - this is accordingly called time-varying confounding. Statistical adjustment for such confounding is quite complicated; Hernán and Robins (2024) devote about a third of their book to it. One of their proposed methods, called the g-formula for time-varying treatments (Section 21.1) could in principle be combined with BART, and some first efforts in this direction have been made by Wang et al. (2024). Extending their work to the CEA setting seems possible in principle, but far beyond the scope of the present paper.
{acks}
[Acknowledgments] No financial support for this study was provided.
Appendix
A Simulation experiment
We demonstrate here, through a simple simulation experiment, the consequences of model misspecification when estimating treatment effects. Such experiments are useful because we know the true data-generating process (hence also the true treatment effects) and can hence investigate how a statistical method performs at recovering the truth. R code to replicate the experiment is given at the end.
The goal of our hypothetical analysis is estimating effect of a treatment $t_{i}$ on the healthcare costs $c_{i}$, and there is a single continuous baseline covariate $x_{i}$. Our true model is given by
$c_{i}=t_{i}+3\sin(x_{i})x_{i}^{2}+\epsilon_{i}$ (3)
In line with common practice, our choice of the statistical model will be a simple linear regression:
$c_{i}=\beta_{0}+\beta_{1}t_{i}+\beta_{2}x_{i}+\epsilon_{i}$ (4)
where $\epsilon_{i}\sim\mathbf{N}(0,\sigma^{2}).$ Under the assumption of correct model specification, we have $\beta_{1}=\Delta_{c}$.
However, our model in (4) is evidently not correctly specified. The question now arises what happens to our estimates, when we fit model (4) regardless of the misspecification. Do they still provide useful information about the treatment effect?
As mentioned in Section 3, the answer depends on whether the treatment assignment depends on the baseline covariate $x_{i}$. We first assign, to each patient, a probability of $\frac{1}{2}$ of receiving treatment $t_{i}=1$, independently of $x_{i}$. The simulation hence mimics a randomized trial.
We now draw $100$ random datasets from our specified simulation, and then analyse each dataset with the linear regression model. The left panel in Figure A.4 shows a histogram of the estimated treatment effects. Despite the misspecification, the estimates concentrate around the true value $1$.
Figure A.4: Estimates of $\Delta_{c}$, obtained through linear regression.
We now modify the simulation such that $t_{i}$ is no longer independent of $x_{i}$, thereby mimicking an observational study, where the treatment assignment depends on the patients’ baseline characteristics.
$\text{Pr}(t_{i}=1)=\frac{1}{1+e^{-3x_{i}}}$ (5)
The probability in (5) is called the propensity score.
The results are shown in the right panel of Figure A.4. Unlike before, we observe a clear bias in the estimates.
While our example was of course rather simplistic, the same patterns can be observed in more complex and realistic simulation experiments. A case in point is the seminal work by Dorie et al. (2019), where a much more elaborate and systematic simulation experiment may be found. The authors simulated data a variety of hypothetical observational studies, using nonlinear data-generating processes. They then had other teams of researchers, who had not seen the true data-generating processes, try to estimate the average treatment effect through a statistical method of their choice. It turned out that all variations of linear regression models performed poorly, as did other kinds of parametric, while some nonparametric regression methods did very well.
To illustrate the utility of nonparametric regression, we rerun the simulation experiment from the previous section, using BART instead of a linear regression. The results are presented in A.5. For each of the two data-generating processes, the inferred treatment effects concentrate around the true value.
Figure A.5: Estimates of $\Delta_{c}$, obtained through BART.
1N <- 200
2X <- rnorm(N)
3ATE_linear_sim1 <- rep(NA, N_sim)
4ATE_linear_sim2 <- rep(NA, N_sim)
5ATE_BART_sim1 <- rep(NA, N_sim)
6ATE_BART_sim2 <- rep(NA, N_sim)
7EY_0 <- 3 * sin(X) * X^2
8EY_1 <- EY_0 + 1
9ATE_true <- mean(EY_1 - EY_0)
10
11N_sim <- 100
12test <- matrix(c(rep(X, 2), rep(0, N), rep(1, N)), ncol = 2)
13
14# sim 1: random assignment
15for (i in 1:N_sim){
16 A <- rbinom(N, 1, 0.5)
17 Y <- (1 - A)*EY_0 + A*EY_1 + rnorm(N)
18 ATE_linear_sim1[i] <- lm(Y ~ A + X)\$coefficients[2]
19 train <- data.frame(X,A)
20 BART_mod <- subart(x_train = train,
21 y_mat = as.matrix(Y),
22 x_test = test,
23 n_tree = 100,
24 n_mcmc = 1100,
25 n_burn = 100,
26 varimportance = FALSE)
27 ATE_BART_sim1[i] <- mean(BART_mod\$y_hat_test_mean[(N+1):(2*N),1] - BART_mod\$y_hat_test_mean[1:N,1])
28}
29
30# sim 2: nonrandom assigment
31for (i in 1:N_sim){
32 A <- rbinom(N, 1, 1/(1 + exp(-3*X)))
33 Y <- (1 - A)*EY_0 + A*EY_1 + rnorm(N)
34 ATE_linear_sim2[i] <- lm(Y ~ A + X)\$coefficients[2]
35 train <- data.frame(X,A)
36 BART_mod <- subart(x_train = train,
37 y_mat = as.matrix(Y),
38 x_test = test,
39 n_tree = 100,
40 n_mcmc = 1100,
41 n_burn = 100,
42 varimportance = FALSE)
43 ATE_BART_sim2[i] <- mean(BART_mod\$y_hat_test_mean[(N+1):(2*N),1] - BART_mod\$y_hat_test_mean[1:N,1])
44}
45
46data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
47 ATE_linear = c(ATE_linear_sim1,ATE_linear_sim2),
48 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
49)
50data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
51 ATE_linear = c(ATE_linear_sim1,ATE_linear_sim2),
52 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
53)
54data_sim\$sim <- factor(data_sim\$sim, levels = c("randomized", "non-randomized"))
55
56ggplot(data_sim) +
57 geom_histogram(aes(x = ATE_linear), color = "black", fill = "grey") +
58 scale_x_continuous(limits = c(-1,3)) +
59 geom_vline(xintercept = 1, color = "red") +
60 facet_wrap(. ~ sim) +
61 theme_classic() +
62 theme(axis.title.x = element_blank())
63
64ggplot(data_sim) +
65 geom_histogram(aes(x = ATE_BART), color = "black", fill = "grey") +
66 scale_x_continuous(limits = c(-1,3)) +
67 geom_vline(xintercept = 1, color = "red") +
68 facet_wrap(. ~ sim) +
69 theme_classic() +
70 theme(axis.title.x = element_blank())
Listing 8: Code for simulation experiment